OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 + ... + 67

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@hartnn

OpenStudy (amistre64):

you need to define the rule, or pattern

OpenStudy (anonymous):

This is the one with the big E in it

OpenStudy (amistre64):

i know, but the meat of it is the rule that defines the sequence

OpenStudy (amistre64):

\[\sum_{n=a}^{b}~f(n)\]

OpenStudy (amistre64):

does your material want you to start with n=0? or n=1? different authors choose different starting indexes

OpenStudy (anonymous):

yeah like that!

OpenStudy (anonymous):

the bottom of the E is n = 0

OpenStudy (amistre64):

you still need to define the rule for f(n)

OpenStudy (anonymous):

idk how to do that :l

OpenStudy (amistre64):

\[\large \sum_{n=0}^{b}~f(n)\]

OpenStudy (anonymous):

Can i show u my the choices? its only 4

OpenStudy (amistre64):

you have a sequence of numbers .... -8, -3, 2, 7, ... , 67 look in your bag of "knowledge" and determine what you see happening

OpenStudy (anonymous):

its increasing by 5 each time. eventually it gets up to 67. first term is -8

OpenStudy (amistre64):

then lets go with a arithmetic sequence rule: -8 +5n , assuming we want n to start at 0 what is the value of n what this thing is equal to 67?

OpenStudy (anonymous):

n = 15

OpenStudy (anonymous):

so 15 on top, n = 0 on bottom

OpenStudy (amistre64):

our rule for f(n) is -8+5n at n=0, -8+5(0) = -8 at n=15, -8+5(15) = 67 \[\large \sum_{n=0}^{15}~-8+5n\] works fine

OpenStudy (anonymous):

yeahhh! thanks amistre!

OpenStudy (amistre64):

yw